OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#17556 — LINX EXTREME 100G, UK, EU
Scheduled Maintenance Report for Network & Infrastructure
Completed
We will add 1 * 100 G with the exchange point LINX EXTREME.
The current port is not configured on a port-channel, we will filter all BGP sessions on this exchange before passing the new 100G production.
This maintenance is scheduled at 01:00 CEST.
Following this maintenance we will have 2 * 100G with the exchange point LINX EXTREME.

Update(s):

Date: 2016-04-28 17:19:41 UTC
After investigation, this time the trouble comes from LINX side configuration.

The port-channel is up for 5 minutes now.

Traffic flows properly and no CRC is detected:

ldn-1-n7# sh inter Eth5/1,Eth6/2 | i \"30 sec\"
Load-Interval #1: 30 seconds
30 seconds input rate 3325256416 bits/sec, 1715106 packets/sec
30 seconds output rate 12524022128 bits/sec, 1533466 packets/sec
Load-Interval #1: 30 seconds
30 seconds input rate 3394268376 bits/sec, 1700938 packets/sec
30 seconds output rate 13435516280 bits/sec, 1622690 packets/sec
ldn-1-n7# sh inter Eth5/1,Eth6/2 | i CRC
1 runts 0 giants 0 CRC/FCS 0 no buffer
0 runts 0 giants 0 CRC/FCS 0 no buffer

Date: 2016-04-28 17:18:24 UTC
Ports are suspended.

Date: 2016-04-28 17:17:24 UTC
We just updated ldn-1-n7 to correct the BUG on LACP: task 17753

3rd try to add the second 100G in the port-channel: We believe in it !! :)

Date: 2016-04-28 17:14:08 UTC
We have reproduced the problem.
The 2 ports are mounted for a few minutes then the LACP negotiation stopped working.
We took the catch packets on the port-channel and other control output for LACP.

We will continue troubleshooting with cisco tomorrow.

Date: 2016-04-28 17:11:53 UTC
Second trial! We have implemented a span to troubleshoot if the problem recurs.

Date: 2016-04-28 17:10:28 UTC
We have detected LACP negotiation problems and ports are suspended after the expiry of LACP Timeout.
LINX will attempt to reproduce the problem in a lab on their side.

We handed traffic 1 * 100 G without using port-channel.

Date: 2016-04-28 17:09:15 UTC
The exchange point is isolated, we will now proceed with the establishment of two 100G ports in the same LAG.

Date: 2016-04-28 17:07:59 UTC
We are isolating the exchange point LINX EXTREME in order to proceed with the upgrade.
Posted Apr 14, 2016 - 17:47 UTC